One earthquake has MMS magnitude 4.4. If a second earthquake has 730 times as much energy (earth movement) as the first, find the magnitude of the second quake.


1
Expert's answer
2021-01-21T03:52:09-0500

For an earthquake with seismic moment S, a measurement of earth movement, the MMS value, or magnitude of the earthquake, is

                                                                               "M = \large\frac{2}{3}log(\large\frac{S}{S_o})"

Where "S_o = 10^{16}" is a baseline measure for the seismic moment.


Let's solve we have following quantities:

            "M_1 = 4.4"           n = 730 = "\large\frac{S_2}{S_1}"                 "M_2 - ?"

"S_1 = S_o 10^{1.5M_1}"

"S_2 = S_o 10^{1.5M_2}"


"\frac{S_2}{S_1} = 10^{1.5(M_2-M_1)} = 730"


"1.5(M_2-M_1) = log730"    "\to M_2 = \frac{2}{3}log730 +M_1 = 6.3"

Answer:   the magnitude of the second quake is 6.3